# Maxitive functions with respect to general orders

### v2

@@ Abstract.
In decision-making, maxitive functions are used for worst-case and best-case evaluations. Maxitivity gives rise to a rich structure that is well-studied in the context of the pointwise order. In this article, we investigate maxitivity with respect to general preorders and provide a representation theorem for such functionals. The results are illustrated for different stochastic orders in the literature, including the usual stochastic order, the increasing convex/concave order, and the dispersive order.
@@
The second author was partially supported by Fundación Séneca - ACyT Región de Murcia project 21955/PI/22, Agencia Estatal de Investigación (Government of Spain) and Project PID2021-122126NB-C32 funded by MICIU/AEI /10.13039/ 501100011033/ and FEDER A way of making Europe.
@@ 1. Introduction
In the context of decision-making under risk and uncertainty, one usually considers a set of gambles $X$ and a functional $\psi\colon X\to\mathbb{R}$, which assigns to each gamble $f\in X$ a number $\psi(f)$ representing its price, risk, or utility. In the theory of imprecise probabilities, these functionals are called previsions, see e.g. [25], while in financial mathematics they are often referred to as risk measures [9] or nonlinear expectations [18].
In many cases, the functional $\psi$ takes the form of an expectation, where the value $\psi(f)$ represents the weighted average of the values associated with the gamble $f$. However, decisions are often based on assessments of best-case or worst-case scenarios, rather than relying on averages. These evaluations are described with maxitive functions. A functional $\psi$ is called maxitive if $\psi(f\vee g)=\psi(f)\vee\psi(g)$ for any two gambles $f$ and $g$, where $f\vee g$ denotes the pointwise supremum of $f$ and $g$.
Maxitive functionals play a crucial role in possibility theory [7, 8, 28] and serve as aggregation functions in several fields such as economics, information fusion, and control theory [10, 24]. Under reasonable assumptions, maxitive functions can be represented in terms of maxitive integrals, such as the Shilkret integral discussed in [23, 2, 19, 13]. Beyond the theory of decision-making, maxitive functions appear in other contexts, including idempotent or tropical analysis [14], or in large deviations theory [12, 13, 20].
In probability theory and statistics, and particularly in decision theory, there exists a wide range of criteria to define order relations between gambles beyond the pointwise order, see [17, 22]. While the majority of the literature has concentrated on maxitivity with respect to pointwise order, little attention has been paid to maxitivity with respect to general orders. The aim of the present work is to close this gap by studying maxitivity with respect to preorders. This allows, in particular, the application of the maxilive structure and its beneficial properties to larger classes of functions. By passing from the pointwise order to weaker stochastic orders, new maxitive functionals are obtained; e.g., it is shown in [26] that the value at risk is maxitive with respect to the usual stochastic order.
@@
The paper is organized as follows: Section 2 covers basic concepts and definitions. In Section 3, we present and prove the main result. Section 4 discusses variations of the main result for different stochastic orders. Finally, in an appendix, we provide representation results for maxitive risk measures on $C_{b}$ and auxiliary results for quantile functions.
@@ 2. Basic concepts and definitions
Let $X$ be a non-empty set and let $\Phi$ be a non-empty set of functions $\phi\colon X\to\mathbb{R}$. Throughout, the set $X$ is endowed with the preorder
$x\preccurlyeq y\quad\mbox{if and only if}\quad\phi(x)\leq\phi(y)\quad\mbox{forall }\phi\in\Phi.$ (2.1)
Further, let $\mathcal{Y}$ be a family of subsets of $X$ which includes all finite subsets. Typical examples of $\mathcal{Y}$ are the family of all finite subsets, the family of all at most countable subsets, or the power set.
@@
$(X,\preccurlyeq)$ is $\mathcal{Y}$-complete if every $Y\in\mathcal{Y}$ which is bounded from above has a unique supremum $\vee Y$.
@@
Let $\psi\colon X\to(-\infty,\infty]$ be a function with sublevel sets $(A_{s})_{s\in\mathbb{R}}$ given by $A_{s}:=\{x\in X\colon\psi(x)\leq s\}$.
  • (1)
    The function $\psi$ is monotone if $x\preccurlyeq y$ implies $\psi(x)\leq\psi(y)$.
  • (2)
    If $(X,\preccurlyeq)$ is $\mathcal{Y}$-complete, then the function $\psi$ is $\mathcal{Y}$-maxitive if for every $Y\in\mathcal{Y}$,
  • (3)
    The function $\psi$ has (upwards) $\mathcal{Y}$-directed sublevel sets if for each $s\in\mathbb{R}$, the set $A_{s}$ is $\mathcal{Y}$-directed, i.e., for every $Y\in\mathcal{Y}$ with $Y\subset A_{s}$, there exists $y^{\ast}\in A_{s}$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$.
These concepts are related as follows.
Lemma 2.3.
Suppose that $(X,\preccurlyeq)$ is $\mathcal{Y}$-complete. Then a function $\psi\colon X\to(-\infty,\infty]$ is $\mathcal{Y}$-maxitive if and only if $\psi$ is monotone and has $\mathcal{Y}$-directed sublevel sets.
@@
First, suppose that $\psi$ is $\mathcal{Y}$-maxitive. Then for $x,y\in X$ with $x\preccurlyeq y$, it follows from $\mathcal{Y}$-maxitivity with $Y:=\{x,y\}$ that $\psi(x)\leq\psi(x\vee y)=\psi(y)$. Moreover, for $Y\in\mathcal{Y}$ with $Y\subset A_{s}$, it holds $\sup_{y\in Y}\psi(y)\leq s$ and $\mathcal{Y}$-maxitivity implies that $Y$ is bounded from above with $\vee Y\in A_{s}$.
Second, if $\psi$ is monotone, then the right hand side of (2.2) is greater than the left hand side. If $\psi$ has $\mathcal{Y}$-directed sublevel sets, then either $\sup_{y\in Y}\psi(y)=\infty$ or $\sup_{y\in Y}\psi(y)=s<\infty$ in which case $\vee Y\in A_{s}$ so that $\psi(\vee Y)\leq s$, i.e., the right hand side of (2.2) is smaller than the left hand side. ∎
As a typical example, we consider the set $X=[-\infty,\infty]^{\Omega}$ of all [,][$\infty$,$\infty$][ - ∞ , ∞ ]-valued functions on a non-empty set $\Omega$, and the family $\Phi:=\{\phi_{\omega}\colon\omega\in\Omega\}$ where $\phi_{\omega}\colon X\to\mathbb{R}$ is given by $\phi_{\omega}(f)=\arctan(f(\omega))$. [^1] We set $\arctan(\infty):=\pi/2$ and $\arctan(-\infty):=-\pi/2$. Then $\Phi$ induces the pointwise order on $X$, and the concept of $\mathcal{Y}$-maxitivity amounts to the usual concepts of maxitivity studied in the literature, see e.g. [2] and references therein. In particular, if $\mathcal{Y}$ is the family of all finite subsets of $X$, then the set of all $\mathcal{Y}$-maxitive functions is exactly the set of all maxitive functions. Likewise, if $\mathcal{Y}$ is the family of countable subsets of $X$ (respectively the family of all subsets of $X$), then the set of all $\mathcal{Y}$-maxitive functions is the set of all countably maxitive functions (respectively the set of all completely maxitive functions).
Definition 2.4.
A function $\alpha\colon{\Phi}\to[-\infty,\infty]$ is $\mathcal{Y}$-stable if for every $Y\in\mathcal{Y}$ with $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, there exists $y^{\ast}\in X$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$ and $\phi(y^{\ast})\leq\alpha(\phi)$ for all $\phi\in\Phi$.
Proposition 2.5.
Suppose that $A=\{x\in X\colon\phi(x)\leq\alpha(\phi)\mbox{ for all }\phi\in\Phi\}$ for some $\alpha\colon\Phi\to[-\infty,\infty]$. Then
$A=\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\},$
where $\alpha_{\min}(\phi):=\sup_{x\in A}\phi(x)$ for all $\phi\in\Phi$. In this case, $A$ is $\mathcal{Y}$-directed, if and only if $\alpha$ is $\mathcal{Y}$-stable, if and only if $\alpha_{\min}$ is $\mathcal{Y}$-stable.
Proof.
First, since $\phi(x)\leq\alpha(\phi)$ for all $x\in A$ and $\phi\in\Phi$, it holds $\alpha_{\min}(\phi)\leq\alpha(\phi)$ for all $\phi\in\Phi$, which shows that $\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\}\subset A$. Conversely, if $x\in A$, it holds $\phi(x)\leq\alpha_{\min}(\phi)$ for all $\phi\in\Phi$, which shows that $A\subset\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\}$.
Second, if $A$ is $\mathcal{Y}$-directed and $Y\in\mathcal{Y}$ satisfies $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, then $Y\subset A$ so that there exists $y^{\ast}\in A$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$, and therefore $\phi(y^{\ast})\leq\alpha(\phi)$ for all $\phi\in\Phi$, i.e., $\alpha$ is $\mathcal{Y}$-stable. Conversely, if $\alpha$ is $\mathcal{Y}$-stable and $Y\in\mathcal{Y}$ satisfies $Y\subset A$, then $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$ and $\mathcal{Y}$-stability implies that $Y$ is bounded from above by some $y^{\ast}\in A$, i.e., $A$ is $\mathcal{Y}$-directed. The argumentation for $\alpha_{\min}$ is analogous. ∎

### v3

@@ Abstract.
In decision-making, maxitive functions are used for worst-case and best-case evaluations. Maxitivity gives rise to a rich structure that is well-studied in the context of the pointwise order. In this article, we investigate maxitivity with respect to general preorders and provide a representation theorem for such functions. The results are illustrated for different stochastic orders in the literature, including the usual stochastic order, the increasing convex/concave order, and the dispersive order.
@@
2010 Mathematics Subject Classification:
We wish to express our gratitude to two anonymous referees for their careful reading of the manuscript and for their comments, which significantly improved the presentation. The second author was partially supported by Fundación Séneca - ACyT Región de Murcia project 21955/PI/22, Agencia Estatal de Investigación (Government of Spain) and Projects PID2021-122126NB-C32 and PID2022-137396NB-I00 funded by MICIU/AEI/10.13039/ 501100011033/ and FEDER A way of making Europe.
@@ 1. Introduction
In decision-making under risk and uncertainty, one typically considers a set of gambles $X$ and a function $\psi\colon X\to\mathbb{R}$, which assigns to each gamble $f\in X$ a value $\psi(f)$, representing its price, risk, or utility. In the context of imprecise probabilities, such functions are referred to as lower/upper previsions; see [31]. In financial mathematics, they are commonly called risk measures [12] or nonlinear expectations [23].
In many cases, the function $\psi$ takes the form of an expectation, where the value $\psi(f)$ represents the weighted average of the values associated with the gamble $f$. However, decisions are often based on assessments of best-case or worst-case scenarios, rather than relying on averages. These evaluations are captured by maxitive functions. A function $\psi$ is called maxitive if it satisfies the condition $\psi(f\vee g)=\max\{\psi(f),\psi(g)\}$ for any two gambles $f$ and $g$, where $f\vee g$ denotes the pointwise supremum of $f$ and $g$.
Maxitive functions play a crucial role in possibility theory [10, 11, 34] and serve as aggregation functions in various fields such as economics, information fusion, and control theory [13, 29]. Under reasonable assumptions, these functions can be represented in terms of maxitive integrals, such as the Shilkret integral discussed in [3, 16, 24, 28]. Beyond decision theory, maxitive functions also arise in other areas, including idempotent or tropical analysis [17] and large deviations theory [15, 16, 25].
In probability theory, statistics, and particularly decision theory, a wide range of criteria exist to define order relations between gambles beyond the pointwise order [22, 27]. While most of the literature has focused on maxitivity with respect to pointwise order, less attention has been paid to maxitivity with respect to more general orders. The aim of this work is to fill this gap by studying maxitivity in the context of preorders, which enables the extension of the maxitive structure and its beneficial properties to a wider class of functions. By passing from the pointwise order to weaker stochastic orders, new maxitive functions are obtained. For example, in [32], it is shown that the value-at-risk is maxitive with respect to the usual stochastic order.
@@
The paper is structured as follows: In Section 2, we provide some motivation for maxitivity in decision-making and possibility theory. Section 3 introduces the basic concepts and definitions. In Section 4, we present and prove the main result. Section 5 explores variations of the main result in the context of different stochastic orders. Finally, the appendix provides representation results for maxitive risk measures on $C_{b}$, along with auxiliary results related to quantile functions.
2. Maxitive functions in decision-making
Let $X$ be a set of gambles of the form $f\colon S\to[-\infty,\infty]$, where each gamble $f$ assigns a loss $f(s)$ to each possible state $s\in S$ [^3] Here, we view negative losses as gains. We note that decision functions are usually defined on spaces of positions. However, for our purposes, describing them in terms of losses seems to be more transparent from a mathematical perspective. Since positions and losses only differ by their sign, all results can be easily transformed. We consider a decision maker who evaluates the gambles using a decision function $\psi\colon X\to(-\infty,\infty]$, which in risk analysis is commonly referred to as a risk measure; see [12]. We adopt the convention that $\psi(f)\leq\psi(g)$ indicates that the gamble $g$ is considered riskier than the gamble $f$. The set of acceptable gambles for the decision maker is defined by the acceptance set, given by $A:=\{f\in X\colon\psi(f)\leq 0\}$.
[…]
$\psi(f\vee g)=\max\{\psi(f),\psi(g)\}\quad\mbox{for all }f,g\in X.$
We also want to mention that in the literature, there are different versions of maxitivity, all of which require $\psi\left(\vee_{i\in I}f_{i}\right)=\sup_{i\in I}\psi(f_{i})$, but for different families $(f_{i})_{i\in I}$, for which the supremum $\vee_{i\in I}f_{i}$ is assumed to exist; see e.g. [3, 28]. For example, $\psi$ is called completely maxitive if this holds for all families $(f_{i})_{i\in I}$, and $\psi$ is called countably maxitive if this holds for countable families $(f_{n})_{n\in\mathbb{N}}$. Furthermore, a topological perspective on maxitivity, which is closely linked to the large deviation principle, was recently investigated in [16]; see also [15, 25].
2.1. Maxitivity for the pointwise order in the context of possibility theory; see [10, 11, 34]
[…]
$\psi(f)=\max_{s\in S}f(s)\pi(s)$
on the space of positive gambles $f\colon S\to[0,\infty]$. We note that this example generalizes to arbitrary state spaces $S$. Moreover, the mapping $\sup_{s\in S}f(s)\pi(s)$ coincides with the Shilkret integral $\int^{S}f\,\mathrm{d}\Pi$ with respect to the possibility measure $\Pi$, which assigns each subset $B\subset S$ the value $\Pi(B):=\sup_{s\in B}\pi(s)$, where $\pi$ is a possibility distribution; see e.g. [3, 5, 19, 28]. Another closely related example is the penalized worst-case evaluation $\sup_{s\in S}\{f(s)-\alpha(s)\}$ for some penalty $\alpha\colon S\to[-\infty,\infty]$. Again this is a completely maxitive function on the space $X$ of all gambles $f\colon S\to[-\infty,\infty]$. Conversely, it is known that every completely maxitive function $\psi\colon X\to[-\infty,\infty]$, which has the translation property and satisfies $\psi(0)\in\mathbb{R}$, admits the representation
$\psi(f)=\sup_{s\in S}\{f(s)-\alpha(s)\}\quad\text{for all }f\in X,$ (2.1)
for some penalty $\alpha\colon S\to[-\infty,\infty]$; see [3, Corollary 7]. The right-hand side in the previous equation corresponds to the so-called convex integral, which is obtained as a transformation of the Shilkret integral; see [3]. In the translation-invariant case (particularly for monetary risk measures), these considerations show that maxitive functions with respect to the pointwise order take the form (2.1) of a penalized worst-case evaluation over the state space.
2.2. Maxitivity with respect to other orders
As an illustration, we start with the usual stochastic order. In this case, we order the distributions of the gambles according to their quantile functions. We recall that for two gambles $f$ and $g$, there exists a gamble $f\vee_{\rm st}g$ whose quantile function is the smallest one that dominates the quantiles $q_{f}$ and $q_{g}$; see [22]. A decision function is then maxitive if $\psi(f\vee_{\rm st}g)=\max\{\psi(f),\psi(g)\}$ for all gambles $f,g$. Intuitively, maxitivity still ensures that losses in certain states cannot be compensated by gains in other states. However, in this context, the relevant ‘states’ refer to the levels of the quantile functions. More formally, suppose we have two gambles $f$ and $g$, where $f$ is acceptable but $f\vee_{\rm st}g$ is not acceptable, for a decision function $\psi$ with acceptance set $A$. If $\psi$ is maxitive, then any gamble $h$ that satisfies $q_{s}(h)=q_{s}(g)$ for all $s$ in the ‘bad states’ $U=\{s\colon q_{g}(s)\geq q_{f}(s)\}$ is not acceptable, regardless of its values $q_{s}(h)$ in the ‘good states’ $s\in U^{c}$. Indeed, since $f\vee_{\rm st}g\preccurlyeq f\vee_{\rm st}h$, it follows that
$0<\psi(f\vee_{\rm st}g)\leq\psi(f\vee_{\rm st}h)=\max\{\psi(f),\psi(h)\}.$
Given that $\psi(f)\leq 0$, this implies $\psi(h)>0$, meaning that $h$ is not acceptable. In other words, maxitivity with respect to the usual stochastic order implies that risky quantiles at certain levels (‘bad states’) cannot be compensated by less risky quantiles at other levels (‘good states’). When dealing with probability distributions, it seems natural to consider quantities (that are closely linked to the distribution) as the relevant ‘states’. A typical example of such a maxitive function is the Value at Risk, which is widely used in the financial industry. Formally, the Value at Risk at level $s\in(0,1)$, defined by ${\rm VaR}_{s}(f):=q_{f}(s)$, is maxitive with respect to the usual stochastic order but is not maxitive with respect to the pointwise order; see [32].
In the context of model aggregation in risk analysis, an interesting motivation for maxitivity with respect to the usual stochastic order was recently presented in [32], where the topic is discussed in great detail. Simplifying their example, consider a risk analyst who evaluates random losses using different methods, resulting in a set of possible risk models $\{f_{1},\dots,f_{n}\}$, which are aggregated into a worst-case model $f^{\ast}:=f_{1}\vee_{\rm st}\cdots\vee_{\rm st}f_{n}$. Under maxitivity, we obtain the desirable property that the risk of the worst-case model $\psi(f^{\ast})$ equals the worst-case risk $\max\{\psi(f_{1}),\dots,\psi(f_{n})\}$. Moreover, it is shown in [32] that every real-valued function $\psi$, which has the translation property, is lower semicontinuous with respect to the convergence in distribution [^5] I.e., $\liminf_{n\to\infty}\psi(f_{n})\geq\psi(f)$ whenever $f_{n}$ converges to $f$ in distribution. and is completely maxitive (with respect the usual stochastic order) admits a representation
$\psi(f)=\sup_{s\in(0,1)}\{q_{f}(s)-\alpha(s)\}\quad\mbox{ for all }f,$ (2.2)
for some penalty $\alpha\colon(0,1)\to[-\infty,\infty]$. The representations in (2.2) and (2.1) are closely related, as (2.2) can be interpreted as an analogue of (2.1), where quantile evaluation replaces scenario-wise evaluation.
Another widely used preorder in decision-making is the dispersive order, which is employed to compare the dispersion or variability of two gambles. Specifically, for two gambles $f,g$, we write $f\preccurlyeq_{\rm disp}g$ if and only if
$q_{f}(v)-q_{f}(u)\leq q_{g}(v)-q_{g}(u)\quad\text{whenever }0<u<v<1.$
For any two gambles $f$ and $g$, there exists a gamble $f\vee_{\rm disp}g$, which is the least upper bound of $f$ and $g$ with respect to $\preccurlyeq_{\rm disp}$. Intuitively, $f\vee_{\rm disp}g$ is the least dispersed gamble that is more dispersed than both $f$ and $g$; see [22]. We say that a function $\psi\colon X\to\mathbb{R}$ is maxitive with respect to $\preccurlyeq_{\rm disp}$ if $\psi(f\vee_{\rm disp}g)=\max\{\psi(f),\psi(g)\}$ for any two gambles $f,g$. In this context, maxitivity once again ensures that losses in certain states cannot be offset by gains in others. Here, the relevant ‘states’ correspond to all pairs of levels $(u,v)$ with $10<u<v<$ 1, at which the quantile spread $q_{g}(v)-q_{g}(u)$ is evaluated. Thereby, maxitivity with respect to the dispersive order implies that risky quantile spreads at certain pairs of levels cannot be compensated by less risky quantile spreads at other levels.
These ideas naturally extend to other partial orders. Consider a decision-maker who compares gambles based on the evaluations of certain relevant features represented by a set $\Phi$ of functions, i.e., $f\preccurlyeq g$ if and only if $\phi(f)\leq\phi(g)$ for all $\phi\in\Phi$ [^6] For the pointwise order, we consider the evaluation $\phi_{s}(f)=f(s)$ for each state $s\in S$. For the usual stochastic order, the evaluation is given by $\phi_{s}(f)=q_{f}(s)$ at each quantile level $s\in(0,1)$. For the dispersive order, we use the evaluation $\phi_{(u,v)}(f)=q_{f}(v)-q_{f}(u)$ for each pair $(u,v)\in\Delta$ of quantile levels. In this situation, assuming that any two gambles have a least upper bound, a decision function $\psi\colon X\to\mathbb{R}$ is called maxitive if it satisfies $\psi(f\vee_{\Phi}g)=\max\{\psi(f),\psi(g)\}$ for all gambles $f,g$. We prove in this paper that, in the translation-invariant case and under reasonable assumptions, maxitive functions can be represented in the form
$\psi(f)=\sup_{\phi\in\Phi}\{\phi(f)-\alpha(\phi)\},$ (2.3)
for some penalty function $\alpha\colon\Phi\to[-\infty,\infty]$. However, in general, a function that admits a representation (2.3) is not necessarily maxitive. To ensure this, the penalty has to satisfy an additional stability condition. Furthermore, we want to emphasize that the $\phi$’s generally do not satisfy the translation property (e.g. for the dispersive order), in which case the representation in (2.3) is more involved.
@@ 3. Basic concepts and definitions
In this section, we provide the formal definitions of maxitivity with respect to general orders and introduce the relevant related concepts. Let $X$ be a non-empty set [^7] In the previous motivation, $X$ was a set of gambles., which is endowed with a preorder
$x\preccurlyeq y\quad\text{if and only if}\quad\phi(x)\leq\phi(y)\quad\text{forall }\phi\in\Phi,$
where $\Phi$ is a non-empty set of functions $\phi\colon X\to\mathbb{R}$. Throughout this article, the notation $Y\subset X$ denotes non-strict inclusion, meaning that every element of $Y$ is also an element of $X$, allowing for the possibility that $Y=X$.
[…]
$\psi(x\vee y)=\max\{\psi(x),\psi(y)\}\quad\mbox{for all }x,y\in X.$
[…]
Lemma 3.1.
Let $(X,\preccurlyeq)$ be an upper semilattice. For a function $\psi\colon X\to(-\infty,\infty]$, the following conditions are equivalent:
  • (i)
    $\psi$ is maxitive.
  • (ii)
    $\psi$ is monotone [^9] I.e., $x\preccurlyeq y$ implies $\psi(x)\leq\psi(y)$. and has upwards directed sublevel sets [^10] I.e., if $x,y\in A_{s}=\{z\in X\colon\psi(z)\leq s\}$ for some $s\in\mathbb{R}$, then there exists $z^{\ast}\in A_{s}$ with $x\preccurlyeq z^{\ast}$ and $y\preccurlyeq z^{\ast}$..
  • (iii)
    $\psi$ is monotone and for every partition $\{\Phi_{1},\Phi_{2}\}$ of $\Phi$ and $x_{1},x_{2}\in X$ with $\phi(x_{1})\geq\phi(x_{2})$ for all $\phi\in\Phi_{1}$ and $\phi(x_{2})\geq\phi(x_{1})$ for all $\phi\in\Phi_{2}$, there exists $i\in\{1,2\}$ such that
    $\phi(y)\geq\phi(x_{i})\mbox{ for all }\phi\in\Phi_{i}\quad\mbox{implies}\quad\psi(y)\geq\psi(x_{1}\vee x_{2}).$
Proof.
${\rm(i)}\Rightarrow{\rm(ii)}$: Given $x,y\in X$ with $x\preccurlyeq y$, we have that $y=x\vee y$. Since $\psi$ is maxitive, we obtain that $\psi(y)=\psi(x\vee y)=\max\{\psi(x),\psi(y)\}\geq\psi(x)$, which shows that $\psi$ is monotone. Moreover, given $x,y\in A_{s}$ for some $s\in\mathbb{R}$, we fix $z^{\ast}=x\vee y$. By maxitivity, it follows that $\psi(z^{\ast})=\psi(x)\vee\psi(y)\leq s$, so that $z^{\ast}\in A_{s}$. This shows that $\psi$ has upwards directed sublevel sets.
${\rm(ii)}\Rightarrow{\rm(i)}$: Given $x,y\in X$, by monotonicity, we have that $\psi(x\vee y)\geq\max\{\psi(x),\psi(y)\}$. Suppose by contradiction that $\psi(x\vee y)>s:=\max\{\psi(x),\psi(y)\}$. Since $x,y\in A_{s}$ and $\psi$ has upwards directed sublevel sets, there exists $z^{\ast}\in A_{s}$ such that $x,y\preccurlyeq z^{\ast}$. Hence, $x\vee y\preccurlyeq z^{\ast}$, which by monotonicity implies that $\psi(x\vee y)\leq\psi(z^{\ast})\leq s$. This contradicts that $\psi(x\vee y)>s$.
$\rm{(i)}\Rightarrow{\rm(iii)}$: By contradiction, assume that there exist a partition $\{\Phi_{1},\Phi_{2}\}$ of $\Phi$ and $x_{1},x_{2}\in X$ with $\phi(x_{1})\geq\phi(x_{2})$ for all $\phi\in\Phi_{1}$ and $\phi(x_{2})\geq\phi(x_{1})$ for all $\phi\in\Phi_{2}$, for which there exist $y_{1},y_{2}\in X$ with
$\phi(y_{i})\geq\phi(x_{i})\mbox{ for all }\phi\in\Phi_{i}\quad\mbox{and}\quad\psi(y_{i})<\psi(x_{1}\vee x_{2}),$
for $i\in\{1,2\}$. Then, for any $\phi\in\Phi$, either $\phi\in\Phi_{1}$ so that $\phi(y_{1}\vee y_{2})\geq\phi(x_{1})\geq\phi(x_{2})$, or $\phi\in\Phi_{2}$ so that $\phi(y_{1}\vee y_{2})\geq\phi(x_{2})\geq\phi(x_{1})$, which shows that $x_{1}\vee x_{2}\preccurlyeq y_{1}\vee y_{2}$. Since $\psi$ is maxitive, it is also monotone, which leads to the following contradiction:
$\psi(x_{1}\vee x_{2})\leq\psi(y_{1}\vee y_{2})=\max\{\psi(y_{1}),\psi(y_{2})\}<\psi(x_{1}\vee x_{2}).$
${\rm(iii)}\Rightarrow{\rm(i)}$: For $x_{1},x_{2}\in X$, we consider the partition $\Phi_{1}:=\{\phi\colon\phi(x_{1})\geq\phi(x_{2})\}$ and $\Phi_{2}:=\{\phi\colon\phi(x_{1})<\phi(x_{2})\}$. Then, there exists $i\in\{1,2\}$ such that
$\phi(y)\geq\phi(x_{i})\mbox{ for all }\phi\in\Phi_{i}\quad\mbox{implies}\quad\psi(y)\geq\psi(x_{1}\vee x_{2}).$
By choosing $y:=x_{i}$ in the previous relation, we obtain that $\psi(x_{1}\vee x_{2})\leq\psi(x_{i})\leq\max\{\psi(x_{1}),\psi(x_{2})\}$. The other inequality $\psi(x_{1}\vee x_{2})\geq\max\{\psi(x_{1}),\psi(x_{2})\}$ follows directly from the monotonicity of $\psi$. ∎
Discussion 3.2.
[…]
Representation results in terms of a weighted or a penalized worst-case evaluation for maxitive functions generally require stronger forms of maxitivity; see e.g. [3, 5]. For instance, $\psi$ is called completely maxitive if
$\psi\big{(}\vee_{i\in I}x_{i}\big{)}=\sup_{i\in I}\psi\left(x_{i}\right)$
for every family $(x_{i})_{i\in I}\subset X$. However, for such a notion to make sense, the supremum $\vee_{i\in I}x_{i}$ must exist for arbitrary families. Under additional regularity assumptions, complete maxitivity is already ensured by (finite) maxitivity [1, 15, 16, 20]. For our purposes, and to achieve the most elegant mathematical representation results, we will equip the ‘state space’ $\Phi$ with a Polish topology and assume that the mapping $\Phi\to\mathbb{R}$, $\phi\mapsto\phi(x)$, is lower semicontinuous for all $x\in X$; see Section 4. In this context, we work with the following version of maxitivity:
@@
Suppose that $(X,\preccurlyeq)$ is countably complete [^13] We say that $(X,\preccurlyeq)$ is countably complete if every countable subset121212 Recall taht we use the convention that $A\subset B$ means that every element of $A$ is an element of $B$. $Y\subset X$ that is bounded from above has a unique supremum $\vee Y$ in $X$. Recall that $Y$ is bounded from above, if there exists $y^{\ast}\in X$ such that $y\preccurlyeq y^{\ast}$ for all $y\in Y$.. Then, a function $\psi\colon X\to(-\infty,\infty]$ is bounded countably maxitive if for every countable subset $Y\subset X$, we have
As in the finite case described in Lemma 3.1, we obtain the following characterization of bounded countable maxitivity:
Lemma 3.4.
Suppose that $(X,\preccurlyeq)$ is countably complete. Then, a function $\psi\colon X\to(-\infty,\infty]$ is bounded countably maxitive if and only if it is monotone and has countably upwards directed sublevel sets [^14] A function $\psi$ has countably upwards directed sublevel sets if for every $s\in\mathbb{R}$, the level set $A_{s}:=\{x\in X:\psi(x)\leq s\}$ is countably upwards directed, i.e., for every countable subset $Y\subset A_{s}$, there exists $y^{\ast}\in A_{s}$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$..
Proof.
First, suppose that $\psi$ is bounded countably maxitive. Then, for $x,y\in X$ with $x\preccurlyeq y$, it follows from bounded countably maxitivity with $Y:=\{x,y\}$ that $\psi(x)\leq\psi(x\vee y)=\psi(y)$. Moreover, for any countable subset $Y\subset A_{s}$, it holds $\sup_{y\in Y}\psi(y)\leq s$ and bounded countable maxitivity implies that $Y$ is bounded from above with $\vee Y\in A_{s}$.
Second, if $\psi$ is monotone, then the right hand side of (3.1) is greater than the left hand side. If $\psi$ has countably upwards directed sublevel sets, then either $\sup_{y\in Y}\psi(y)=\infty$ or $\sup_{y\in Y}\psi(y)=s<\infty$ in which case $\vee Y\in A_{s}$ so that $\psi(\vee Y)\leq s$, i.e., the right hand side of (3.1) is smaller than the left hand side. ∎
Our representation results are based on duality arguments. Hereby, maxitivity translates dually into the following stability property.
@@
A function $\alpha\colon{\Phi}\to[-\infty,\infty]$ is countably stable if for every countable subset $Y\subset X$ with $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, there exists $y^{\ast}\in X$ with $y\preccurlyeq y^{\ast}$ for all $y\in Y$ and $\phi(y^{\ast})\leq\alpha(\phi)$ for all $\phi\in\Phi$.
At the level of sublevel sets, we have the following characterization.
Proposition 3.6.
Suppose that
$A=\{x\in X\colon\phi(x)\leq\alpha(\phi)\text{ for all }\phi\in\Phi\},$
for some $\alpha\colon\Phi\to[-\infty,\infty]$. Then, we have
$A=\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\text{ for all }\phi\in\Phi\},$
where $\alpha_{\min}(\phi):=\sup_{x\in A}\phi(x)$ for all $\phi\in\Phi$.
In this case, $A$ is countably upwards directed if and only if $\alpha$ is countably stable if and only if $\alpha_{\min}$ is countably stable.
@@
First, since $\phi(x)\leq\alpha(\phi)$ for all $x\in A$ and $\phi\in\Phi$, it holds that $\alpha_{\min}(\phi)\leq\alpha(\phi)$ for all $\phi\in\Phi$, which shows that $\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\}\subset A$. Conversely, if $x\in A$, then $\phi(x)\leq\alpha_{\min}(\phi)$ for all $\phi\in\Phi$, which shows that $A\subset\{x\in X\colon\phi(x)\leq\alpha_{\min}(\phi)\mbox{ for all }\phi\in\Phi\}$.
Second, if $A$ is countably upwards directed and the subset $Y\subset X$ is countable with $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, then $Y\subset A$, so there exists $y^{*}\in A$ with $y\preccurlyeq y^{*}$ for all $y\in Y$, and therefore $\phi(y^{*})\leq\alpha(\phi)$ for all $\phi\in\Phi$, i.e., $\alpha$ is countably stable. Conversely, if $\alpha$ is countably stable and $Y\subset X$ is countable such that $Y\subset A$, then $\sup_{y\in Y}\phi(y)\leq\alpha(\phi)$ for all $\phi\in\Phi$, and countable stability implies that $Y$ is bounded from above by some $y^{*}\in A$, i.e., $A$ is countably upwards directed. The argument for $\alpha_{\min}$ is analogous. ∎